# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ Appendix A Solution to Remark 1
Combining equations (B)–(B) and substituting $\mathbf{U}_{i}=c\mathbf{B}_{i}^{T}\mathbf{B}_{i}+c\mathbf{A}_{i}^{T}\mathbf{A}_{i}+\mathbf{H}_{i}^{T}\mathbf{H}_{i}$, we derive:
@@
$+$ $\mathbf{y}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\big((\mathbf{q}_{i,j}^{-})^{t+1}-(\mathbf{q}_{i,j}^{-})^{t}\big)$
$+$ $\mathbf{y}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\big((\mathbf{q}_{i,j}^{+})^{t+1}-(\mathbf{q}_{i,j}^{+})^{t}\big),\,j\in\mathcal{N}_{i}\Big).$
Let $\mathbf{1}_{3N_{i}n}\in\mathbb{R}^{N_{i}n}$ denote a $3N_{i}n$-dimensional column vector of all ones. Then it follows:
$\mathbf{1}_{3N_{i}n}^{T}\mathbf{A}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})=\sum_{j\in\mathcal{N}_{i}}\Big[$ $\mathbf{x}_{i}^{t+1}-\mathbf{x}_{i}^{t}-((\mathbf{p}_{i,j}^{-})^{t+1}-(\mathbf{p}_{i,j}^{-})^{t})$
$+$ $\mathbf{y}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\big((\mathbf{q}_{i,j}^{-})^{t+1}-(\mathbf{q}_{i,j}^{-})^{t}\big)$
$+$ $\mathbf{y}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\big((\mathbf{q}_{i,j}^{+})^{t+1}-(\mathbf{q}_{i,j}^{+})^{t}\big)\Big].$ (B.24)
Substituting equations (34a)–(34f) into equation (B) and summing over nodes $i\in\mathcal{N}$ yields:
$\sum_{i\in\mathcal{N}}\mathbf{1}_{3N_{i}n}^{T}\mathbf{A}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})=\sum_{i\in\mathcal{N}}\sum_{j\in\mathcal{N}_{i}}$ $\Big[\frac{c}{c+1}\Big(\tilde{\mathbf{x}}_{i}^{t+1}-\mathbf{x}_{i}^{t}-\Big((\tilde{\mathbf{p}}_{i,j}^{-})^{t+1}-(\mathbf{p}_{i,j}^{-})^{t}\Big)$
$+\tilde{\mathbf{y}}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\Big((\tilde{\mathbf{q}}_{i,j}^{-})^{t+1}-(\mathbf{q}_{i,j}^{-})^{t}\Big)$
$+\tilde{\mathbf{y}}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\Big((\tilde{\mathbf{q}}_{i,j}^{+})^{t+1}-(\mathbf{q}_{i,j}^{+})^{t}\Big)$
@@
For the first three lines on the right-hand side of equation (B), using equation (19) again, we derive:
@@
$+\tilde{\mathbf{y}}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\Big((\tilde{\mathbf{q}}_{i,j}^{-})^{t+1}-(\mathbf{q}_{i,j}^{-})^{t}\Big)$
$+\tilde{\mathbf{y}}_{i}^{t+1}-\mathbf{y}_{i}^{t}-\Big((\tilde{\mathbf{q}}_{i,j}^{+})^{t+1}-(\mathbf{q}_{i,j}^{+})^{t}\Big)\Big]$
@@
where $\mathbf{C}_{i}=[c\cdot\mathbf{I}_{N_{i}n},(c+1)\cdot\mathbf{I}_{N_{i}n},(c+1)\cdot\mathbf{I}_{N_{i}n}]\in\mathbb{R}^{n\cdot N_{i}\times n\cdot 3N_{i}}$.
For the fourth line on the right-hand side of equation (B), since $\mathbf{z}_{i}^{t}$ is the optimal solution to problem (31), we have $\mathbf{z}_{i}^{t}\in\mathcal{Y}$ for all $t\geq 1$, and thus:
@@
Combining equation (B.27), the fourth line on the right-hand side of equation (B) can be written as:
@@
$=$ $\sum_{i\in\mathcal{N}}\frac{1}{c+1}\mathbf{1}_{N_{i}n}^{T}\mathbf{M}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}).$ (B.28)
where $\mathbf{M}_{i}=[\mathbf{I}_{N_{i}n},\mathbf{0}_{N_{i}n}]\mathbf{H}_{i}\in\mathbb{R}^{n\cdot N_{i}\times n\cdot(N_{i}+1)}$. The last equality holds based on the expressions for matrix $\mathbf{H}_{i}$ in equation (II-C) and the definitions of vectors $\mathbf{z}_{i}^{t+1},\mathbf{z}_{i}^{t}$.
Substituting equations (B) and (B) into equation (B), we obtain:
$\left\|\sum_{i\in\mathcal{N}}\mathbf{1}_{3N_{i}n}^{T}\mathbf{A}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})\right\|^{2}=\frac{1}{(c+1)^{2}}\left\|\sum_{i\in\mathcal{N}}\mathbf{1}_{N_{i}n}^{T}\left[\mathbf{C}_{i}\mathbf{A}_{i}+\mathbf{M}_{i}\mathbf{H}_{i}\right](\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})\right\|^{2}.$ (B.29)
@@
$\|\mathbf{1}_{3N_{\text{sum}}n}^{T}\mathbf{A}(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}=\frac{1}{(c+1)^{2}}\|\mathbf{1}_{N_{sum}n}^{T}[\mathbf{CA}+\mathbf{MH}](\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}.$ (B.30)
@@
$3N_{\text{sum}}n\|\mathbf{A}(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}\geq\|\mathbf{1}_{3N_{\text{sum}}n}^{T}A(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}.$ (B.31)
@@
where $\tilde{\tau}_{\min}$ is defined as the smallest nonzero eigenvalue of $\mathbf{1}_{N_{sum}n}^{T}(\mathbf{CA}+\mathbf{MH})]^{T}[\mathbf{1}_{N_{sum}n}^{T}(\mathbf{CA}+\mathbf{MH})]$. Based on the definitions of $\mathbf{A}_{i},\mathbf{H}_{i}$ in equations (II-C) and (II-C), it can be derived that:
$\tilde{\tau}_{\min}=\min\quad\{5(c+1)^{2}N_{i}^{2}+(3c^{2}+4c+3)N_{i},i\in\mathcal{N}\}.$
@@
$\frac{c}{2}\|\mathbf{A}(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}\geq\frac{c\cdot\tilde{\tau}_{\min}}{6N_{sum}n(c+1)^{2}}\|\tilde{\mathbf{z}}^{t+1}-\mathbf{z}^{t}\|^{2}.$
@@
$\sum_{i\in\mathcal{N}}\frac{c}{2}\|\mathbf{A}_{i}(\mathbf{z}^{t+1}_{i}-\mathbf{z}_{i}^{t})\|^{2}\geq\sum_{i\in\mathcal{N}}\frac{c\cdot\tilde{\tau}_{\min}}{6N_{sum}n(c+1)^{2}}\|\tilde{\mathbf{z}}^{t+1}_{i}-\mathbf{z}^{t}_{i}\|^{2}.$ (B.34)
Finally, substituting equation (B.34) into equation (B) completes the proof.
The result of Lemma 1 follows from Lemmas 4, 5 and 6, along with the potential function $\mathcal{P}^{t}$ defined in Equation (IV).
Appendix C (Proof of Theorem 1)
@@
The update step of $\bm{\lambda}_{i}^{t+1}$ in equation (26c), together with equations (B.8), (B), and (C.1), leads to
@@
Combining the above with equation (D.1), we get that $\mathbf{z}_{i}^{t}$ is the optimal solution of problem (D). Therefore, using the first-order optimality condition of problem (D), we obtain
@@
$\mathbf{w}_{i}^{t}=\mathrm{proj}_{\mathcal{B}^{N_{i}+1}}\left(\mathbf{w}_{i}^{t-1}-\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})\right),$
@@
$0=\|\mathbf{w}_{i}^{t-1}-\mathbf{w}_{i}^{t}\|^{2}=\|\mathbf{w}_{i}^{t-1}-\mathrm{proj}_{\mathcal{B}^{N_{i}+1}}\left(\mathbf{w}_{i}^{t-1}-\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})\right)\|^{2}.$
@@
$\min_{\mathbf{w}_{i}\in\mathcal{B}^{N_{i}+1}}\|\mathbf{w}_{i}-\mathbf{w}_{i}^{t-1}+\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})\|^{2}.$ (D.7)
@@
$0\leq\langle\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t}),\mathbf{w}_{i}-\mathbf{w}_{i}^{t}+\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1}\rangle=\langle\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t}),\mathbf{w}_{i}-\mathbf{w}_{i}^{t}\rangle,\quad\forall\mathbf{w}_{i}\in\mathcal{B}^{N_{i}+1}.$ (D.8)
@@
$\mathbf{w}_{i}^{t}\in\arg\min_{\mathbf{w}_{i}\in\mathcal{B}^{N_{i}+1}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t}),$ (D.9)
@@
$0\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})+\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i}^{t}).$ (D.10)
@@
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}\right)+\mathbf{x}_{i}^{t}-\mathbf{y}_{1}^{t}-r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in(\mathcal{N}/\mathcal{A}),$ (D.11)
$\mathbf{0}$ $=\sum_{i\in\mathcal{N}/\mathcal{A}}(\mathbf{y}_{i}^{t}-\mathbf{x}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t})+\sum_{i\in\mathcal{A}}(\mathbf{y}_{i}^{t}-\mathbf{a}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t}),$ (D.12)
$\mathbf{x}_{k}^{t}$ $=\mathbf{a}_{k},\quad\forall k\in\mathcal{A}.$ (D.13)
@@
$=-\text{vec}\big(-d_{i,j}(\mathbf{x}_{i}^{t}-(\mathbf{z}_{i,j}^{+})^{t}),\quad j\in\mathcal{N}_{i}\quad,-r_{i}(\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t})\big)$
$=-\text{vec}\big(-d_{i,j}(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}),\quad j\in\mathcal{N}_{i}\quad,-r_{i}(\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t})\big),$ (D.15)
where
$\mathbf{D}_{i}$ $=\operatorname{Diag}\left(\operatorname{vec}(d_{i,j},j\in\mathcal{N}_{i}),r_{i}\right)\otimes\mathbf{I}_{n},\,\mathbf{H}_{i}=\begin{bmatrix}\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}.$
$\mathbf{z}_{i}$ $=[\mathbf{x}_{i},vec(\mathbf{p}_{i,j}^{-},j\in\mathcal{N}_{i})^{T},vec(\mathbf{p}_{i,j}^{+},j\in\mathcal{N}_{i})^{T},\mathbf{y}_{i},vec(\mathbf{q}_{i,j}^{-},j\in\mathcal{N}_{i})^{T},vec(\mathbf{q}_{i,j}^{+},j\in\mathcal{N}_{i})^{T}]^{T},\forall i\in\mathcal{N}.$ (D.16)
@@
$G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})$ $=\frac{1}{2}\|\mathbf{H}_{i}\mathbf{z}_{i}\|^{2}-\mathbf{w}_{i}^{T}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i},\quad$
$\mathcal{X}:$ $=\{\mathbf{z}\mid\mathcal{E}_{i}\mathbf{z}_{i}=\mathbf{a}_{i},\,\forall\,i\in\mathcal{A}\},$ (D.17)
$\text{where}\quad\mathcal{E}_{i}:$ $=\begin{bmatrix}1,&\mathbf{0}_{4N_{i}+1}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\times(2+4N_{i})n}.$
Substituting equations (D.16) and (D) into (43a), the optimization problem (43a) is rewritten as

### v2

@@ A.1 Solution to Remark 2
Combining equations (A.2)–(A.2) and substituting $\mathbf{U}_{i}=c\mathbf{B}_{i}^{T}\mathbf{B}_{i}+c\mathbf{A}_{i}^{T}\mathbf{A}_{i}+\mathbf{H}_{i}^{T}\mathbf{H}_{i}$, we derive:
@@
$+$ $\sum_{k\in\mathcal{M}}\big(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t}\big)$
$+$ $\sum_{k\in\mathcal{M}}\big(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t}\big)\big)\big),\,j\in\mathcal{N}_{i}\Big).$
Let $\mathbf{1}_{(1+2M)N_{i}n}\in\mathbb{R}^{N_{i}n}$ denote a $(1+2M)N_{i}n$-dimensional column vector of all ones. Then it follows:
$\mathbf{1}_{(1+2M)N_{i}n}^{T}\mathbf{A}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})=\sum_{j\in\mathcal{N}_{i}}\Big[$ $\mathbf{x}_{i}^{t+1}-\mathbf{x}_{i}^{t}-((\mathbf{p}_{i,j}^{-})^{t+1}-(\mathbf{p}_{i,j}^{-})^{t})$
$+$ $\sum_{k\in\mathcal{M}}\Big(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t}$
$+$ $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t}\big)\big)\Big)\Big].$ (B.24)
Substituting equations (33a)–(33f) into equation (A.2) and summing over nodes $i\in\mathcal{N}$ yields:
$\sum_{i\in\mathcal{N}}\mathbf{1}_{(1+2M)N_{i}n}^{T}\mathbf{A}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})=\sum_{i\in\mathcal{N}}\sum_{j\in\mathcal{N}_{i}}$ $\Big[\frac{c}{c+1}\big(\tilde{\mathbf{x}}_{i}^{t+1}-\mathbf{x}_{i}^{t}-\big((\tilde{\mathbf{p}}_{i,j}^{-})^{t+1}-(\mathbf{p}_{i,j}^{-})^{t}\big)$
$+\sum_{k\in\mathcal{M}}\Big(\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t}\big)$
$+\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t}\big)\Big)$
@@
For the first three lines on the right-hand side of equation (A.2), using equation (18) again, we derive:
@@
$+\sum_{k\in\mathcal{M}}\Big(\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t}\big)$
$+\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}^{t+1}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\big((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t}\big)\Big)\Big]$
@@
where $\mathbf{C}_{i}=[c\cdot\mathbf{I}_{N_{i}n},(c+1)\cdot\mathbf{I}_{N_{i}n},...,(c+1)\cdot\mathbf{I}_{N_{i}n}]\in\mathbb{R}^{n\cdot N_{i}\times n\cdot(1+2M)N_{i}}$.
For the fourth line on the right-hand side of equation (A.2), since $\mathbf{z}_{i}^{t}$ is the optimal solution to problem (30), we have $\mathbf{z}_{i}^{t}\in\mathcal{Y}$ for all $t\geq 1$, and thus:
@@
Combining equation (B.27), the fourth line on the right-hand side of equation (A.2) can be written as:
@@
$=$ $\sum_{i\in\mathcal{N}}\frac{1}{c+1}\mathbf{1}_{N_{i}n}^{T}\mathbf{M}_{i}\mathbf{H}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}).$ (B.28)
where $\mathbf{M}_{i}=[\mathbf{I}_{N_{i}n},\mathbf{0}_{N_{i}n},...,\mathbf{0}_{N_{i}n}]\in\mathbb{R}^{n\cdot N_{i}\times n\cdot(N_{i}+M)}$. The last equality holds based on the expressions for matrix $\mathbf{H}_{i}$ in equation (19) and the definitions of vectors $\mathbf{z}_{i}^{t+1},\mathbf{z}_{i}^{t}$.
Substituting equations (A.2) and (A.2) into equation (A.2), we obtain:
$\left\|\sum_{i\in\mathcal{N}}\mathbf{1}_{(1+2M)N_{i}n}^{T}\mathbf{A}_{i}(\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t})\right\|^{2}=\frac{1}{(c+1)^{2}}\left\|\sum_{i\in\mathcal{N}}\mathbf{1}_{N_{i}n}^{T}\left[\mathbf{C}_{i}\mathbf{A}_{i}+\mathbf{M}_{i}\mathbf{H}_{i}\right](\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})\right\|^{2}.$ (B.29)
@@
$\|\mathbf{1}_{(1+2M)N_{\text{sum}}n}^{T}\mathbf{A}(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}=\frac{1}{(c+1)^{2}}\|\mathbf{1}_{N_{sum}n}^{T}[\mathbf{CA}+\mathbf{MH}](\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}.$ (B.30)
@@
$3N_{\text{sum}}n\|\mathbf{A}(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}\geq\|\mathbf{1}_{(1+2M)N_{\text{sum}}n}^{T}A(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}.$ (B.31)
@@
where $\tilde{\tau}_{\min}$ is defined as the smallest nonzero eigenvalue of $\mathbf{1}_{N_{sum}n}^{T}(\mathbf{CA}+\mathbf{MH})]^{T}[\mathbf{1}_{N_{sum}n}^{T}(\mathbf{CA}+\mathbf{MH})]$. Based on the definitions of $\mathbf{A}_{i},\mathbf{H}_{i}$ in equations (19) and (17), it can be derived that:
$\tilde{\tau}_{\min}=\min\quad\{(c+1)^{2}[(6M+1)N_{i}^{2}+N_{i}],i\in\mathcal{N}\}.$
@@
$\frac{c}{2}\|\mathbf{A}(\mathbf{z}^{t+1}-\mathbf{z}^{t})\|^{2}\geq\frac{c\cdot\tilde{\tau}_{\min}}{2(2M+1)N_{sum}n(c+1)^{2}}\|\tilde{\mathbf{z}}^{t+1}-\mathbf{z}^{t}\|^{2}.$
@@
$\sum_{i\in\mathcal{N}}\frac{c}{2}\|\mathbf{A}_{i}(\mathbf{z}^{t+1}_{i}-\mathbf{z}_{i}^{t})\|^{2}\geq\sum_{i\in\mathcal{N}}\frac{c\cdot\tilde{\tau}_{\min}}{2(2M+1)N_{\text{sum}}n(c+1)^{2}}\|\tilde{\mathbf{z}}^{t+1}_{i}-\mathbf{z}^{t}_{i}\|^{2}.$ (B.34)
Finally, substituting equation (B.34) into equation (A.2) completes the proof.
The result of Lemma 1 follows from Lemmas 4, 5 and 6, along with the potential function $\mathcal{P}^{t}$ defined in Equation (4).
A.3 (Proof of Theorem 1)
@@
The update step of $\bm{\lambda}_{i}^{t+1}$ in equation (25c), together with equations (B.8), (A.2), and (C.1), leads to
@@
Combining the above with equation (D.1), we get that $\mathbf{z}_{i}^{t}$ is the optimal solution of problem (A.4). Therefore, using the first-order optimality condition of problem (A.4), we obtain
@@
$\mathbf{w}_{i}^{t}=\mathrm{proj}_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}\left(\mathbf{w}_{i}^{t-1}-\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})\right),$
@@
$0=\|\mathbf{w}_{i}^{t-1}-\mathbf{w}_{i}^{t}\|^{2}=\|\mathbf{w}_{i}^{t-1}-\mathrm{proj}_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}\left(\mathbf{w}_{i}^{t-1}-\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})\right)\|^{2}.$
@@
$\min_{\mathbf{w}_{i}\in\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}\|\mathbf{w}_{i}-\mathbf{w}_{i}^{t-1}+\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})\|^{2}.$ (D.7)
@@
$0\leq\langle\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t}),\mathbf{w}_{i}-\mathbf{w}_{i}^{t}+\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1}\rangle=\langle\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t}),\mathbf{w}_{i}-\mathbf{w}_{i}^{t}\rangle,\quad\forall\mathbf{w}_{i}\in\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}.$ (D.8)
@@
$\mathbf{w}_{i}^{t}\in\arg\min_{\mathbf{w}_{i}\in\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t}),$ (D.9)
@@
$0\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i}^{t}).$ (D.10)
@@
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}\right)+\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}}(\mathbf{x}_{i}^{t}-\mathbf{y}_{1,k}^{t}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t}),\quad\forall i\in(\mathcal{N}/\mathcal{A}),$ (D.11)
$\mathbf{0}$ $=\sum_{i\in\mathcal{N}/\mathcal{A}}\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{x}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t})+\sum_{i\in\mathcal{A}}\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{a}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,k}^{t}),$ (D.12)
$\mathbf{x}_{l}^{t}$ $=\mathbf{a}_{l},\quad\forall l\in\mathcal{A}.$ (D.13)
@@
$=-\text{vec}\big(-d_{i,j}(\mathbf{x}_{i}^{t}-(\mathbf{z}_{i,j}^{+})^{t}),\quad j\in\mathcal{N}_{i}\quad,-r_{i,{\color[rgb]{0,0,0}k}}(\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}),\quad k\in\mathcal{M}\big)$
$=-\text{vec}\big(-d_{i,j}(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}),\quad j\in\mathcal{N}_{i}\quad,-r_{i,{\color[rgb]{0,0,0}k}}(\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}),\quad k\in\mathcal{M}\big),$ (D.15)
@@
$G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})$ $=\frac{1}{2}\|\mathbf{H}_{i}\mathbf{z}_{i}\|^{2}-\mathbf{w}_{i}^{T}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i},\quad\mathcal{X}:=\{\mathbf{z}\mid\mathbf{E}_{i}\mathbf{z}_{i}=\mathbf{a}_{i},\,\forall\,i\in\mathcal{A}\},$ (D.17)
where $\mathbf{E}_{i}:=\begin{bmatrix}1,&\mathbf{0}_{(2N_{i}+1)M+2N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\times(2N_{i}+1)(1+M)n}.$
Substituting equations (D.17) into (42a), the optimization problem (42a) is rewritten as